### v1

# PixWizard: Versatile Image-to-Image Visual Assistant with Open-Language Instructions

Abstract
This paper presents a versatile image-to-image visual assistant, PixWizard, designed for image generation, manipulation, and translation based on free-from language instructions. To this end, we tackle a variety of vision tasks into a unified image-text-to-image generation framework and curate an Omni Pixel-to-Pixel Instruction-Tuning Dataset. By constructing detailed instruction templates in natural language, we comprehensively include a large set of diverse vision tasks such as text-to-image generation, image restoration, image grounding, dense image prediction, image editing, controllable generation, inpainting/outpainting, and more. Furthermore, we adopt Diffusion Transformers (DiT) as our foundation model and extend its capabilities with a flexible any resolution mechanism, enabling the model to dynamically process images based on the aspect ratio of the input, closely aligning with human perceptual processes. The model also incorporates structure-aware and semantic-aware guidance to facilitate effective fusion of information from the input image. Our experiments demonstrate that PixWizard not only shows impressive generative and understanding abilities for images with diverse resolutions but also exhibits promising generalization capabilities with unseen tasks and human instructions. The code and related resources are available at https://github.com/AFeng-x/PixWizard
1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Touvron et al., 2023) and Large Vision Models (LVMs) (Radford et al., 2021; Caron et al., 2021; Bai et al., 2024) have gained global popularity by successfully unifying multiple tasks within a single, coherent framework. Nowadays, LLMs have become efficient language assistants, demonstrating strong capabilities in open-world language understanding and reasoning. However, a versatile visual assistant capable of following diverse multimodal instructions that align with human intentions to effectively perform various visual tasks in real-world scenarios is still under exploration.
Recently, there are two research lines aiming to achieve general visual assistants: diffusion-based and in-context learning approaches. The first focuses on developing text-to-image models (Rombach et al., 2022b) as a unified foundation model for various visual perception tasks. For example, InstructPix2Pix (Brooks et al., 2023), InstructDiffusion (Geng et al., 2024), and InstructCV (Gan et al., 2024) repurpose generative models as a universal language interface for image editing and other visual tasks. However, their capable tasks are limited, and their performance lags behind that of task-specific models. The second research direction focuses on visual prompting, where pixel-based prompts are used to tackle various vision tasks within a single learning framework. This approach employs prompting techniques to generate the desired visual outputs in an in-context manner. Examples include Painter (Wang et al., 2023b), PromptDiffusion (Wang et al., 2023c), and LVM (Bai et al., 2024), which have successfully handled a variety of visual scenarios within one framework. However, these methods inherently lack the ability to follow human language instructions, limiting their controllability and interactivity. (More related works are discussed in Sec. A.)
Figure 1: Task Overview of our Omni Pixel-to-Pixel Instruction-tuning Dataset for PixWizard.
Taking into account the strengths and limitations of previous approaches, we introduce PixWizard, a versatile interactive image-to-image visual assistant designed for image generation, manipulation, and image-to-image translation. PixWizard is a DiT-based framework that can handle a wide range of visual tasks when provided with sufficient training data, along with the interface for human language instructions. Specifically, PixWizard exhibits three main features as follows:
  1. 1.
    Task Unification: Given the diverse nature of vision tasks and data formats, ranging from pixels and coordinates to binary masks and categories, it is challenging to find a unified representation. We observe that most vision tasks can be framed as image-to-image translation problems. For tasks not naturally suited to image output, we first learn to generate their visualizations and then apply post-processing to convert them into the desired formats. This approach marks an initial step toward building a versatile visual assistant.
  2. 2.
    Data Construction: We aim to leverage and integrate the remarkable diversity of tasks and data in the visual domain. To achieve this, we have built a comprehensive training set with a total of 30 million data points, enabling our model to support five main capabilities: (i) Image generation, which includes text-to-image generation, controllable generation, inpainting, and outpainting; (ii) Image editing; (iii) Image restoration, covering tasks such as deraining, desnowing, deblurring, super-resolution, and more; (iv) Image grounding, which involves locating objects based on user prompts; and (v) Dense image prediction, which includes depth estimation, surface normal estimation, pose estimation, semantic segmentation, and image-to-canny/HED/sketch/Line-Art conversions.
  3. 3.
    Architecture Design: For a robust visual assistant, the architecture and scalability of the foundational model are crucial. In this work, we use the flow-based Diffusion Transformer (DiT) (Ma et al., 2024) as our base model. Compared to traditional diffusion models, flow-based models offer greater versatility and stability in modeling data distributions by learning conditional velocity fields. The DiT architecture further enhances scalability and is well-suited for incorporating conditional information. Building on this foundation, we introduce several unique and effective components. Specifically, we extend the dynamic partitioning and padding scheme to enable the model to handle input images of any resolution, closely aligning with human perceptual processes. Additionally, we introduce structure-aware and semantic-aware guidance, allowing the model to effectively follow multimodal instructions (images and open-ended user prompts) for a variety of manipulations.
Experimental results show that PixWizard delivers competitive performance compared to task-specific vision models and surpasses current state-of-the-art general visual models. More importantly, our model handles tasks and instruction prompts it has not encountered during training, demonstrating impressive generalization capabilities. This further highlights PixWizard’s strength as a powerful interactive image-to-image visual assistant.
2 Omni Pixel-to-Pixel Instruction-Tuning Dataset
To empower our image-to-image visual assistant with the capabilities on comprehensive image generation, manipulation, and translation, we first compile a multi-task, multi-model training dataset for visual instruction tuning, consisting of 30 million instances across seven primary domains. To our knowledge, this dataset is the largest, most diverse, and easiest-to-use image-instruction-image triplet dataset. It is constructed from both open-source and in-house datasets, filtered with the assistance of MLLMs and manual review, as illustrated in Fig. 1.
Image Restoration. We incorporate low-level data to restore images degraded by various environmental or technical factors. This section utilizes a wide range of open-source datasets covering key restoration tasks, including (1) Denoising, (2) Deraining, (3) Demoireing, (4) Dehazing, (5) Deblurring, (6) Desnowing, (7) Deshadowing, (8) Low-Light Enhancement, (9) Flash Reflection Removal, (10) Face Restoration, (11) Watermark Removal, and (12) Super Resolution. Since both the inputs and outputs are inherently defined in the RGB space, these tasks can be seamlessly unified by our PixWizard model without extra any transformations. All open-source datasets we use are provided in Sec. B.1.
Image Grounding. Image grounding involves identifying and highlighting specific areas of objects in images based on provided text prompts. The data for this part is sourced from well-known datasets such as gRefCOCO (Liu et al., 2023a), RefCOCO3 (Yu et al., 2016), and Visual Genome (Krishna et al., 2017). We focus on three types of grounding tasks: (1) Segmentation Referring, where the target object specified by the user is highlighted in the output image; (2) Box Detection Referring, where the target object is highlighted using bounding boxes; and (3) Binary Mask Prediction, where a binary mask is directly predicted. (Details are provided in the Sec. B.2.)
Controllable Generation. Referring to ControlNet (Zhang et al., 2023a), we aim to equip our model with natural image generation capabilities given conditional inputs. We first gather natural images from the LAION Art dataset (Schuhmann et al., 2022) and our own collection of high-quality images from the Internet. We then use MiniCPM-Llama3-V 2.5 (Yao et al., 2024), a robust edge-side multimodal LLM, along with advanced techniques to generate captions and conditional inputs for the images. (Details on data construction can be found in Sec. B.3.)
Dense Image Prediction. Dense image prediction tasks require the model to interpret input images and produce dense annotations. For depth estimation, surface normal estimation, and semantic segmentation, we represent per-pixel labels as RGB images, which can be generated via the image generation capabilities. For other tasks, such as the prediction of human pose maps, sketches, HED boundaries, canny edge maps, and cartoon line art, we treat them as image-to-image translation tasks, as we can easily obtain image pairs using open-source tools. (Details are provided in the Sec. B.4.)
Image Editing. Instruction-based image editing holds great potential for practical applications, as it allows users to perform edits using natural language commands. To enhance our model’s ability to modify images according to specific user instructions, we unify several public editing datasets, including UltraEdit (2024), MagicBrush (2024a), GQA-Inpaint (2023), Instruct P2P (2023), SEED-X-Edit (2024), GIER (2020), and HQ-Edit (2024). These datasets encompass a wide range of semantic entities, varying levels of detail, and multiple editing tasks including object removal, object replacement, object addition, background replacement, and style transfer.
Inpainting, in the context of image generation, involves filling in missing parts of an image with new or modified content. To create inpainting instances, we apply random black or white masks to different regions of the original images. These masks come in various shapes, including circles, rectangles, and free-form patterns, and are randomly placed on the images, resulting in a wide range of occluded areas for inpainting.
Outpainting (Image extrapolation) extends an image’s content beyond its boundaries. To create outpainting instances, we randomly crop the central part of the image, while masking the surrounding areas with black or white. The cropping is not limited to square shapes but includes rectangles of varying proportions. This approach ensures that the outpainting task covers a range of extension scenarios, challenging the model to generate coherent content beyond the original image boundaries.
Text-to-Image Generation. Image-to-image and text-to-image are two distinct tasks. The former requires an additional input image as a condition, while the latter does not. Existing instruction-tuning methods transform pretrained text-to-image models into image-to-image systems, but they fail to retain their original text-to-image capabilities. To unify both tasks within a single framework, we propose a "drawing" strategy to preserve the text-to-image capabilities. Specifically, we introduce an additional input alongside the language instructions: an image that is either entirely white or black. This setup simulates a blank canvas, allowing the model to "draw" the generated images based on the provided text prompts. This approach distinguishes our model from previous text-to-image systems. All the image-text pairs we use are sourced from our own collection of high-quality images from the Internet, with captions generated using MLLMs assistance.
Open-language Instruction. To enhance the usability of PixWizard as a practical visual assistant, we aim for the model to understand free-form user prompts. Instead of relying on fixed task-specific prompts, we begin by manually writing 6-10 prompts for each vision task to describe the task. We then use GPT-4o to generate a wide range of paraphrased variations. This process ensures that our instruction set remains diverse while staying true to the core intent of the original prompts. Instruction templates and examples are provided in Sec. B.5.
Figure 2: Overall framework of PixWizard.
3 PixWizard
PixWizard is a versatile image-to-image model fine-tuned on our custom-built omni pixel-to-pixel instruction-tuning dataset. In this section, we present the details of the PixWizard framework from the perspectives of model architecture (as shown in Fig. 2) and training strategies.
3.1 Flow-based Conditional Instruction-Tuning
Previous studies (Wang et al., 2022b; Brooks et al., 2023) show that fine-tuning large diffusion models outperforms training models from scratch for image translation and editing tasks. Therefore, we initialize the weights of PixWizard with the pretrained Lumina-Next-T2I (Zhuo et al., 2024) checkpoint, which is a flow-based diffusion transformer, to leverage its extensive text-to-image generation capabilities. We learn a network $v_{\theta}$ that predicts the velocity field $u_{t}$ given image conditioning $c_{I}$ and text instruction conditioning $c_{T}$. We minimize the loss function as follow:
$\mathcal{L}=\mathbb{E}_{t,p_{1}(x_{1}),p_{t}(x_{t}|x_{1}),c_{I},c_{T}}\|v_{\theta}(x_{t},t,c_{I},c_{T})-u_{t}(x_{t},t|x_{1})\|^{2}.$ (1)
3.2 Architecture
Text Encoders. We begin by using Gemma-2B (Team et al., 2024) as the text embedder in PixWizard to encode text prompts. However, in multi-task learning, relying solely on text instructions is insufficient to guide the model in accurately executing user commands. To better guide the generation process for the correct task, we incorporate the CLIP text encoder (Radford et al., 2021). Global average pooling is applied to the CLIP text embeddings to obtain a coarse-grained text representation, which is passed through an MLP-based task embedder to generate the task embedding. This embedding is then integrated into the PixWizard Block by adding it to the timestep embeddings through a modulation mechanism. As shown in Fig. 3, this approach adaptively clusters similar task instructions in the latent space while separating those from different tasks, which helps guide the model’s generation process in the correct task direction.
Figure 3: t-SNE visualization of the global text embeddings. Each dot represents a human instruction.
Structural-Aware Guidance. To effectively capture the overall structural features of the input image condition, we begin by encoding the image using a variational autoencoder (VAE) (Kingma & Welling, 2013) from SDXL (Podell et al., 2023). Next, we concatenate the image latent with the noise latent along the channel dimension. Following (Brooks et al., 2023), additional input channels are added to the patch embedder, with the weights for these new channels initially set to zero.
Figure 4: The schematic illustrations of PixWizard Block and Task-Aware Dynamic Sampler.
Semantic-Aware Guidance. Besides recognizing structural features, semantic information is also crucial. Therefore, we use CLIP L/14-336 (Radford et al., 2021) to obtain semantic image embeddings. We introduce two zero-initialized attention mechanism within the PixWizard block, which allows the latent target image tokens to query information from the keys and values of the conditions. Specifically, we employ a zero-initialized gating mechanism to gradually inject conditional image and text information into the token sequences. Given target image queries $Q_{i}$, keys $K_{i}$, and values $V_{i}$, along with text instruction keys $K_{t}$ and values $V_{t}$, and conditional image keys $K_{ci}$ and values $V_{ci}$, the final attention output is formulated as:
$A=\text{softmax}\left(\frac{\tilde{Q}_{i}\tilde{K}_{i}^{T}}{\sqrt{d}}\right)V_{i}+\text{tanh}(\alpha_{t})\,\text{softmax}\left(\frac{\tilde{Q}_{i}K_{t}^{T}}{\sqrt{d}}\right)V_{t}+\text{tanh}(\alpha_{ci})\,\text{softmax}\left(\frac{\tilde{Q}_{i}K_{ci}^{T}}{\sqrt{d}}\right)V_{ci},$ (2)
where $\tilde{Q}_{i}$ and $\tilde{K}_{i}$ stand for applying RoPE (Su et al., 2024), $d$ is the dimension of queries and keys, and $\alpha$ indicates the zero-initialized learnable parameter in gated cross-attention.
However, inputting all image tokens into the attention layer often leads to significant computational demands, and we also found that not all semantic tokens are relevant to the specific task. To address this, we introduce the Task-Aware Dynamic Sampler, which is designed to select the most relevant semantic tokens for each task. This sampler uses a lightweight ranking network consisting of four linear layers and activation functions. Inspired by DynamicViT (Rao et al., 2021), we employ a technique that maps image tokens to both local and global features. Furthermore, we integrate task embeddings ${x_{task}}$ to help the sampler identify the tokens that are most relevant to the task. The computational process is formulated as:
$z=MLP(x+x_{task})\in\mathbb{R}^{N\times C},$ (3)
${z}^{\text{local}}=z[:,:\frac{C}{2}]\in\mathbb{R}^{N\times\frac{C}{2}},{z}^{\text{global}}=Avg({z}[:,\frac{C}{2}:])\in\mathbb{R}^{1\times\frac{C}{2}},{z^{\prime}}_{i}=[{z}_{i}^{\rm local},{z}^{\rm global}],\quad 1\leq i\leq N,$ (4)
${M}=MLP({z^{\prime}})\in\mathbb{R}^{N\times 1}$ (5)
where ${M}_{i}$ denotes the importance of the $i$-th token. However, implementing token sparsification is challenging in practice. Directly sampling tokens based on their importance scores is non-differentiable, which hinders end-to-end training. To address this, we use the Gumbel-Softmax technique (Jang et al., 2016) and adapt it into a Multi-Hot Gumbel-Softmax (MHGS) to enable simultaneous sampling of the top $K$ tokens:
$\hat{{x}}={x}\odot MHGS({M})$ (6)
where the output of GumbelSoftmax is a multi-hot tensor that represents the mask of the retained tokens. $\odot$ is the Hadamard product, indicating that the top $K$ tokens by importance score are weighted by 1 and thus retained, while the remaining $(N-K)$ tokens are weighted by zero and discarded. Finally, we equip each layer of the transformer block with an independent task-aware dynamic sampler. This approach not only helps capture the most relevant semantic features needed by each layer to fulfill the requirements of different tasks but also reduces the computational cost in the attention process.
3.3 Any Resolution
PixWizard inherits the dynamic partitioning and padding scheme introduced by (Zhuo et al., 2024), enabling the model to handle images of any resolution and aspect ratio during both fine-tuning and inference. However, in practice, the required resolutions for different tasks can vary significantly. To support more flexible handling of arbitrary resolutions and to preserve the original resolution of each image as much as possible, we use $[512^{2},768^{2},1024^{2}]$ as resolution centers to generate a set of candidate patch partitions. During training, we group data items with similar resolutions into the same bucket to ensure that the longest and shortest sequences within each batch are not too different in length, thereby minimizing the use of padding tokens and improving training efficiency. During inference, by incorporating NTK-Aware Scaled RoPE (Peng & Quesnelle, 2023) and sandwich normalization, PixWizard also demonstrates exceptional resolution extrapolation.
3.4 Two-Stage Training and Data Balancing Strategies
To unlock the model’s potential and improve its performance on tasks with smaller datasets, we propose a two-stage training and data-balancing strategy. S1: At this stage, we initialize the model’s weights by combining the weights of a pretrained text-to-image model with randomly initialized weights for the newly added modules. We begin by selecting tasks with smaller datasets and assigning each dataset a sampling weight to increase its data volume. This weight determines how many times the dataset is repeated during a single epoch. Using this method, each task has approximately 20k data points. We then randomly select training samples from other tasks to match this scale, creating our first-stage training dataset. The training process lasts for a total of 4 epochs. S2: In the second stage, we initialize the model using the weights obtained from the first stage and combine all collected data for further training. To balance tasks, we manually assign sampling weights to each dataset, randomly selecting data if a weight is less than 1.0. We also include text-to-image data at a 1:1 ratio with other tasks, resulting in our second-stage training dataset. At this stage, the total amount of training data reaches up to 20 million samples.
4 Experiments
4.1 First Section Results
Table 1: Comparison of PixWizard with task-specific and vision generalist baselines across six representative tasks, covering both high-level visual understanding and low-level image processing. ’$\times$’ indicates that the method is incapable of performing the task.
Depth Est. Semantic Seg. Surface Normal Est. Denoise Derain Image Grounding
Methods RMSE↓ mIoU↑ Mean Angle Error↓ PSNR↑ SSIM↑ PSNR↑ SSIM↑ cIoU↑ (val set)
NYUv2 SUNRGB-D ADE20K NYU-Depth V2 SIDD Rain100L RefCOCO RefCOCO+
DepthAnything (2024a) 0.206 -
Marigold (2024a) 0.224 -
Mask DINO (2023) 60.80
Mask2Former (2022) 56.10
Bae et al. (2021) 14.90
InvPT (2022) 19.04
AirNet (2022) 38.32 0.945 32.98 0.951
PromptIR (2024) 39.52 0.954 36.37 0.972
LAVT (2022) 72.73 56.86
ReLA (2023a) 73.21 56.10
Unified-IO (2022b) 0.387 0.287 25.71 - $\times$ $\times$ $\times$ $\times$ 46.42 40.50
Painter (2023b) 0.288 0.285 49.90 $\times$ 38.71 0.954 29.87 0.882 $\times$ $\times$
InstructCV (2024) 0.297 0.279 47.23 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
InstructDiffusion (2024) $\times$ $\times$ $\times$ $\times$ 34.26 0.938 19.82 0.741 41.64 33.20
PixWizard 0.287 0.288 36.76 19.65 38.67 0.957 31.43 0.917 46.44 36.49
Settings. For image restoration, we follow previous works (Conde et al., 2024; Potlapalli et al., 2024) and prepare datasets for various restoration tasks during training. For evaluation, we first select two representative benchmarks: Rain100L (2017) for deraining and SIDD (2018) for denoising. Additionally, to further assess performance on other restoration tasks and evaluate zero-shot capabilities, more details and results are provided in the Sec. D.1.
For image grounding, we evaluate referring segmentation tasks on the gRefCOCO (2023a), RefCOCO, and RefCOCO+ validation and test sets. To assess the performance gap with specialized models, we report results from several expert methods and primarily compare our approach with two unified models: Unified-IO and InstructDiffusion. Following standard practices (Liu et al., 2023a), we use cumulative IoU (cIoU) as the performance metric.
Dense image prediction tasks are evaluated on the three vision tasks: ADE20k (2017b) for semantic segmentation, NYUv2 (2012) and SUNRGB-D (2014) for monocular depth estimation, and NYU-Depth v2 (2012) fir surface normal estimation. For semantic segmentation, we assign labels by identifying the nearest neighbor RGB color value, and accuracy is evaluated using the Mean Intersection over Union (mIoU) metric. For monocular depth estimation, we average the output image across the three channels and apply the inverse of the linear transformation used during training to obtain depth estimates within the range of $[0,10]$ meters. Accuracy is evaluated using the Root Mean Square Error (RMSE). For surface normal estimation, we recover the corresponding normal vectors from the output image and use the Mean Angle Error to assess accuracy.
Results. Table 1 presents a comprehensive performance comparison with recent state-of-the-art (SOTA) task-specific and all-in-one methods. As shown in the results, despite denoising and deraining data making up only a small portion of the overall training set, our method outperforms other unified methods and even surpasses some task-specific approaches. In the image grounding task, PixWizard significantly outperforms the diffusion-based generalist model InstructDiffusion by 4.8 cIoU on RefCOCO (val). However, there is still room for improvement compared to other highly specialized models. Furthermore, as shown in Fig.6, PixWizard supports flexible instructions, allowing it to not only highlight and visualize the target object directly on the image but also generate the corresponding binary mask. This highlights its strong performance in real-world interactions and practical applications. Additional quantitative evaluation results are provided in the Sec. D.2.
Figure 5: Qualitative Evaluation of Instruction-based Image Restoration.
Figure 6: Qualitative Results of Instruction-based Image Grounding.
For dense prediction tasks, PixWizard performs competitively against both generalist and task-specific baselines across all three tasks. In depth estimation on the NYUv2 test set, PixWizard achieves a 10.0% improvement in RMSE compared to Unified-IO and performs similarly to Painter and InstructCV. For semantic segmentation, PixWizard surpasses Unified-IO by +11.05 in mIoU, though it still trails behind other methods. Additionally, Fig. 7 provides examples of PixWizard’s outputs. As shown, by supplying the corresponding task-specific prompt for the same image, we can easily generate the respective condition visualization, underscoring PixWizard’s significant practical value.
Figure 7: Visualizations of dense image prediction examples.
4.2 Second Section Results (Image Editing)
Settings. We evaluate PixWizard across two benchmarks, the MagicBrush Test (Zhang et al., 2024a) and the Emu Edit Test (Sheynin et al., 2024), to assess its effectiveness in image editing capabilities. For a fair comparison, we primarily compare it with instruction-guided image editing methods, including InstructPix2Pix, MagicBrush, Emu Edit, and UltraEdit. Consistent with Emu Edit, we use L1 distance, CLIP image similarity, DINO similarity, CLIP text-image similarity, and CLIP text-image direction similarity as metrics.
Table 2: Comparison with image-editing baselines evaluated on Emu Edit and MagicBrush test set.
Emu Edit Test set MagicBrush Test Set
[…]
InstructPix2Pix (2023) 0.078 0.834 0.219 0.121 0.762 0.115 0.837 0.245 0.093 0.767
MagicBrush (2024a) 0.090 0.838 0.222 0.100 0.776 0.123 0.883 0.261 0.058 0.871
Emu Edit (2024) 0.109 0.859 0.231 0.094 0.819 0.135 0.897 0.261 0.052 0.879
UltraEdit (2024) 0.107 0.844 0.283 0.071 0.793 - 0.868 - 0.088 0.792
PixWizard 0.104 0.845 0.248 0.069 0.798 0.124 0.884 0.265 0.063 0.876
Figure 8: Qualitative examples comparing PixWizard with other editing approaches.
Results. Table 2 presents our results compared to the baselines. The findings show that our model consistently outperforms InstructPix2Pix, MagicBrush, and UltraEdit in automatic metrics and achieves comparable performance to state-of-the-art method Emu Edit. Qualitative comparisons are provided in Fig. 8. Our model precisely identifies the editing region while preserving the rest of the pixels, and it demonstrates the best understanding of the given instructions.
4.3 Third Section Results (Image Generation)
Settings. In this section, we focus on evaluating the effectiveness of PixWizard’s generation capabilities. Specifically, we assess its performance across four key tasks: classic text-to-image generation, controllable image generation, inpainting, and outpainting.
For controllable image generation, we evaluated PixWizard’s ability to generate images based on two specific conditions: canny edge maps and depth maps. Following ControlNet++ (Li et al., 2024), we assessed controllability by measuring the similarity between the input conditions and the corresponding features extracted from the generated images. Specifically, we used RMSE for depth map control and F1-Score for canny edges. Additionally, to evaluate the quality of the generated images and their alignment with the input text, we reported FID (Fréchet Inception Distance) and CLIP-Score metrics. All experiments were conducted at a resolution of $512\times 512$.
For image inpainting—where masked regions of an image are filled with new content—we use latent diffusion settings (Rombach et al., 2022b) to measure FID and LPIPS, which assess the quality of the generated samples when 40-50% of the image area needs to be inpainted. For image extrapolation (outpainting), we follow MaskGIT (Chang et al., 2022a) settings, extending the image by 50% to the right and comparing performance against common baselines using FID and Inception Score (IS). Both tasks are evaluated on 30,000 image crops, each sized $512\times 512$, from the Places dataset (Zhou et al., 2017a).
For the text-to-image generation task, we use two primary evaluation methods. First, we visually present examples of images generated by PixWizard. Additionally, we calculate two automatic evaluation metrics: the Human Preference Score (HPS) v2 (Wu et al., 2023) and the standard Zero-shot FID-30K on the MS-COCO dataset (Lin et al., 2014).
Table 3: Comparison of PixWizard with task-specific and vision generalist baselines across six representative tasks, covering both high-level visual understanding and low-level image processing. ’$\times$’ indicates that the method is incapable of performing the task
Canny-to-Image Depth-to-Image Inpainting Outpainting Text-to-Image
Methods FI↑ FID↓ CLIP-S↑ RMSE↓ FID↓ CLIP-S↑ FID↓ LPIPS↓ FID↓ IS↑ FID↓ HPSv2↑
MultiGen-20M MultiGen-20M Places Places COCO-30K Photo
ControlNet-SD1.5 (2023a) 34.65 14.73 32.15 35.90 17.76 32.45
T2I-Adapter-SD1.5 (2024) 23.65 15.96 31.71 48.40 22.52 31.46
LDM-4 (2022b) 9.39 0.246
LaMa (2022) 12.0 0.24
DeepFill v2 (2019) 11.51 17.70
MaskGIT (2022b) 7.80 22.95
DALL·E 2 (2021) 10.32 27.24 ± 0.198
SD 1.5 (2022b) 9.62 27.46 ± 0.198
PixArt-$\alpha$ (2024b) 7.32 -
Lumina-Next (2024) 9.79 27.47 ± 0.203
PixWizard 35.46 15.76 32.01 33.83 16.94 31.84 9.27 0.25 7.54 22.18 9.56 27.47 ± 0.183
Controllable Generation Results. Without the need for separate training for each model, PixWizard is an all-in-one solution capable of handling multiple conditions. As shown in Table 3, PixWizard achieves the highest controllability and best image quality under depth conditions, while also being comparable to current separate models in image-text alignment. Fig. 9 presents several visual samples, demonstrating the effectiveness of our approach.
Figure 9: Visualization examples under different conditions.
Inpainting Results. The comparison with other inpainting approaches in Table 3 shows that PixWizard improves overall image quality, as indicated by FID and LPIPS. We attribute this to PixWizard’s ability to "paint" on a blank canvas, which greatly enhances its capacity to identify masked areas and generate coherent content within them. With this inpainting capability, PixWizard allows for more precise image editing tasks: (i) Remove Anything tackles the object removal problem (Criminisi et al., 2003; 2004; Elharrouss et al., 2020), enabling users to remove specific objects while keeping the result visually seamless. The process consists of two steps: identifying and removing. As shown in Fig. 10, the user first instructs PixWizard on which object to locate, and PixWizard generates a target mask. The mask is then applied to the original image, and PixWizard fills the area with appropriate background details. (ii) Replace Anything lets users swap out any object in an image. The process is similar to Remove Anything, but instead of just removing the object, the model replaces it with the specified object while ensuring the background remains visually consistent. (iii) Add Anything enables users to insert any object into an image and place it wherever they like. The user adds a mask to the desired area and provides a text prompt specifying what to add. PixWizard then leverages its powerful inpainting ability to generate the requested content.
Figure 10: Visualization results of Remove, Replace and Add Anything.
Outpainting Results. As shown in the quantitative comparison results in Table 3, PixWizard outperforms other baselines in the outpainting task, delivering state-of-the-art image generation quality with a FID score of 7.54 and an IS score of 22.18. The samples in Fig. 11 demonstrate PixWizard’s ability to synthesize images in various scenes and styles, and it flexibly handles image extrapolation in multiple directions and aspect ratios with better marginal coherence.
Figure 11: Visualization results of Outpainting.
Text-to-Image Results. As shown in the quantitative comparison results in Table 3, PixWizard achieves a FID score of 9.56 when tested for zero-shot performance on the COCO dataset. Although some models achieve a lower FID, they focus solely on text-to-image tasks and rely on significantly more training resources. Additionally, we evaluated the Human Preference Score (HPS v2), a robust benchmark for assessing human preferences in text-to-image synthesis. PixWizard performed well, delivering image quality comparable to popular text-to-image generators. We also provide visual samples in Fig. 12. PixWizard supports high-resolution image synthesis, up to $1024\times 1024$, with any resolution and aspect ratio.
Figure 12: Text-to-image samples generated by PixWizard.
4.4 Ablation Study
In the Sec. D.3, we conduct ablation studies to examine how the design of each component and the training process impact learning.
5 Discussion and Conclusion
In this work, we explore how to build a versatile interactive image-to-image visual assistant from three key aspects: task definition, data construction, and model architecture. Our goal is to create a system that can precisely follow free-form user instructions for image generation, manipulation, and translation. Our approach, PixWizard, eliminates the need for task-specific design choices and achieves highly competitive performance across a diverse set of tasks, with strong generalization capabilities.
However, this work has some limitations. First, the current model architecture does not yet support multi-image input conditions, which is an increasingly important and valuable research area. Second, there is still room for improvement, particularly in challenging tasks such as segmentation and image grounding when compared to specialized models. Additionally, the performance of the text encoder and foundation model plays a crucial role. Better text encoding methods enable the model to more accurately understand and execute human instructions, while larger and more robust model architectures directly improve the quality of the final output. Notably, the modules and strategies we propose in PixWizard can be easily applied to other powerful text-to-image generators. In the future, we will explore using more advanced diffusion foundation models, such as SD3 and FLUX, and continue to push this promising direction forward until we achieve the vision field’s "GPT-4 moment."
Appendix A Related Work
Diffusion Models. Diffusion models estimate the data distribution by modeling the gradient of the noise-perturbed data distributions (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Dhariwal & Nichol, 2021; Song et al., 2021). They have demonstrated remarkable performance in various fields, ranging from text-to-image generation (Rombach et al., 2022a; Saharia et al., 2022; Betker et al., 2023; Jiang et al., 2024a; Zhang et al., 2023b), controllable generation (Zhang et al., 2023a; Ye et al., 2023), and image editing (Avrahami et al., 2022; Brooks et al., 2023; Kawar et al., 2023) to video (Ho et al., 2022; Brooks et al., 2024), audio (Kong et al., 2021; Huang et al., 2023), 3D (Guo et al., 2023; 2024), and motion (Tevet et al., 2023; Zhang et al., 2024b) generation. Beyond generation, recent works have also exhibited diffusion models’ capabilities in computer vision tasks, such as semantic segmentation (Baranchuk et al., 2022; Xu et al., 2023), depth estimation (Ke et al., 2024b; Lee et al., 2024), and image restoration (Xia et al., 2023). Benefiting from the visual knowledge learned from large-scale pretraining, these works open up the potential for adapting pretrained diffusion models to downstream tasks in a generative manner, with excellent capabilities like handling inherent uncertainties and zero-shot generalization.
Vision Generalists. Building a vision generalist capable of unifying various visual tasks has been a long-standing goal. Inspired by the success of scaling sequential modeling with transformers in natural language processing, many vision generalists (Wang et al., 2022a; Chen et al., 2022b; Lu et al., 2022b; 2024; Mizrahi et al., 2024; Bachmann et al., 2024; Bai et al., 2024) follow a similar approach by converting inputs and outputs into sequences of discrete tokens (Van Den Oord et al., 2017; Esser et al., 2021), allowing joint modeling of different modalities within a unified framework. With the rapid advancements in large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Zhang et al., 2024c) and multi-modal large language models (MLLMs) (Gao et al., 2024; Zhang et al., 2024e; d; Jiang et al., 2024b), several works (Koh et al., 2024; Dong et al., 2024; Sun et al., 2023; Wu et al., ; Han et al., 2023) have introduced task-specific tokens, aligning them with the embedding space of LLMs to equip text-only LLMs with the ability to perceive and generate images. However, a common drawback of these approaches is their limited performance on generation tasks, such as text-to-image generation and image editing. Additionally, their sampling efficiency is constrained by the next-token prediction paradigm, which worsens for image-to-image tasks or when working with high-resolution images. In contrast, diffusion models, known for their state-of-the-art generation performance and efficiency, are ideal candidates for image generation and manipulation tasks. Recent efforts (Wang et al., 2023c; Gan et al., 2024; Hu et al., 2024; Geng et al., 2024) aim to create a visual generalist by unifying multiple visual tasks through a natural language interface based on pretrained text-to-image diffusion. However, these models generally focus on a limited set of tasks within narrow domains and are constrained by the scalability limitations of early foundation models (Rombach et al., 2022b), limiting their potential as practical visual assistants. To address these challenges, we propose a solution from both the model and data perspectives to develop a more versatile visual assistant.
Appendix B More details for the Omni pexel-to-pexel instruction-tuning dataset
B.1 Image Restoration
All the open-source datasets used for the image restoration task are listed below.
BSD (2001) RealBlur (2020) DPDD (2020) GoPro (2017) REDS (2019)
DenseHaze (2019) NH-HAZE (2020) Reside-6K (2018) UHDM (2022) SIDD (2018)
Rain1400 (2017) Outdoor-Rain (2019b) SSID (2022) RainDrop (2018) RainDS (2021)
SRD (2017) RealSnow (2023) Snow100K (2018) CLWD (2021) CelebA-HQ (2015)
RealSR (2019) LOL-v2 (2021) DIV2K (2017) FFHQ (2019) Flickr2K (2023a)
B.2 Image Grounding
(1) Segmentation Referring. We define referring segmentation as highlighting the target object specified by the user in the output image. For example, if the model is given instructions like, "Please mark the pixels in {color} based on the referring description: {caption}," the resulting image would display a mask in the specified color over the appropriate object described in the caption. When constructing the data, we pre-set the mask to a specific color with 50% opacity and apply it directly to the original image. This method makes it easier for humans to evaluate the accuracy of the predicted mask.
(2) Box Detection Referring. Instead of pixel-level grounding, we use bounding boxes to highlight the target object specified by the user in the output image. Prompts for this task include instructions like, "Mark the specified area with a bounding box in {color}: {caption}." The model then frames the described object with a bounding box in the specified color. Similar to referring segmentation, we pre-set the bounding box to a specific color and apply it directly to the original image to produce the final output. During inference, we follow the post-processing methods outlined in InstructCV (Sec. A.3) (Gan et al., 2024) to derive the coordinates of the specified region from the output image.
(3) Binary Mask Prediction. To promote the use of referring segmentation in real-world scenarios, we shift the objective from rendering images to directly predicting a binary mask. The prompt for this task might be: "Generate a binary mask for the described object: {caption}." The model is expected to produce a binary mask image where the object described in the caption is represented as a white region, with the background in black, excluding any original image content. Since the binary mask is a single-channel image, we replicate the mask across three channels to convert it into RGB space during training.
B.3 controllable generation
Canny Edge to Image. We use a Canny edge detector (Canny, 1986) (with random thresholds) and a Tiny and Efficient Edge Detector (TEED) (Soria et al., 2023) to obtain 1M canny-image-caption pairs from our collected natural images.
Holistically-Nested Edge(HED) Boundary to Image. We use HED boundary detection (Xie & Tu, 2015) to obtain 1M edge-image-caption pairs from our collected natural images (a part of images are source of the Canny Edge dataset.)
Depth Map to Image. Depth information is crucial for producing images with a sense of three-dimensionality. We used the Depth Anything V2 model (Yang et al., 2024b) obtain 1M depth-image-caption pairs, enabling accurate generation of depth maps across different visual scenarios.
User Sketch to Image. Following ControlNet (Zhang et al., 2023a), we generate human sketches from images by applying HED boundary detection (Xie & Tu, 2015) combined with strong data augmentations, including random thresholds, random masking of sketch portions, random morphological transformations, and random non-maximum suppression. This process results in 0.5 million sketch-image-caption pairs.
Human Pose to Image. For human pose-based generation, we employed the OpenPose model (Cao et al., 2017) for real-time multi-person 2D pose estimation. To ensure quality, only images where at least 30% of the key points of the whole body were detected were retained; those with fewer detected key points were discarded. We directly use visualized pose images with human skeletons as input condition. Finally, we obtain 0.4M pose-image-caption pairs.
Semantic Segmentation to Image. The semantic mask annotation is produced using OneFormer (Jain et al., 2023), as adopted in ControlNet-1.1  [^1] https://github.com/lllyasviel/ControlNet-v1-1-nightly, providing precise segmentation maps as conditions for image generation. This process results in 1M seg-image-caption pairs.
Normal Map to Image. The surface normals are generated using DSINE Bae & Davison (2024), which contributed to the depiction of surface orientation and texture details in the generated images. Finally, we obtain 0.8M normal-image-caption pairs.
Line-art to Image. We use a cartoon line drawing extraction method (Xiang et al., 2021) to generate line drawings from cartoons. This process yields 0.8M normal-image-caption pairs.
B.4 Dense Image Prediction
Depth Estimation. Monocular depth estimation is a dense prediction task to estimate the per-pixel depth value(distance relative to the camera) given an input RGB image. The datasets we use include Hypersim (Roberts et al., 2021) and VKITTI 2 (Cabon et al., 2020), and our collected depth maps from controllable dataset. For different datasets, the ground-truth depth for each pixel may vary, being either an absolute depth value or a relative depth value. Here we uniformly map the ground-truth value from real-valued range to the integer RGB space with range $[0,255]$, and let the three channels be the same ground truth. During inference, we directly average the outputs of the three channels, and then perform an inverse linear transformation specific to each dataset to obtain a depth estimate in the real-valued range.
Surface Normal Estimation. For surface normal estimation, we train the model to generate RGB images that encode pixel orientations differently. We convert the $x/y/z$ orientations into $r/g/b$ values to create a normal map visualization. The datasets we use include NYUv2 (Silberman et al., 2012) and our collected surface normals maps from controllable dataset.
Semantic Segmentation. Semantic segmentation is a dense prediction task to predict the per-pixel semantic label given an input image. Given a semantic segmentation task, we formulate different semantic categories using different colors in RGB space. Specifically, we define the background and ignore areas as black, i.e., pixels in color $(0,0,0)$, and generate the colors for foreground categories using the pre-defined color-label dictionary. During inference, to decode the output image to a single channel map where each pixel represents a class label, we compute the L2 distance between each output pixel and the pre-defined colors for each semantic category, and take the label of the closest color as the predicted category. The datasets we use is ADE20K (Zhou et al., 2017b), which covers a broad range of 150 semantic categories.
B.5 Instruction Prompt Examples.
Handling free-form user prompts, rather than fixed task-specific prompts, increases PixWizard’s usability as a general visual assistant. In Fig. 13, we present examples of task templates rephrased by GPT-4o.
Figure 13: Examples of GPT4o-paraphrase user prompts for different task.
Appendix C More Detailed Descriptions for the Architecture
C.1 Preliminaries
Diffusion Transformers. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Dhariwal & Nichol, 2021; Song et al., 2021) are a family of generative models demonstrating remarkable performance in modeling data distributions. These models are trained to estimate clean data samples (or added noise) from their noisy version perturbed by a pre-defined Gaussian noise schedule. During sampling, they can generate data samples by iterative denoising starting from prior Gaussian distributions. Recent advancements in diffusion models, such as SoRA (Brooks et al., 2024), SD3 (Esser et al., 2024), and PixArt (Chen et al., 2024b; a), exhibiting a paradigm shift from the classic diffusion U-Net architecture to diffusion transformers (DiTs) (Peebles & Xie, 2023). DiTs appear to be a unified architecture with minimal modifications to original transformers (Vaswani et al., 2017). Therefore, these models demonstrate better scaling properties and can be naturally extended to more modalities, such as integrating text and image conditions through cross-attention.
Flow-based Models. Another line of generative models that extends the definition of diffusion models is ODE-based continuous normalizing flows (Chen et al., 2018), which are also known as flow matching (Lipman et al., 2023; Tong et al., 2024; Albergo & Vanden-Eijnden, 2023) or rectified flows (Liu et al., 2023b). Specifically, given the data space $\mathbb{R}^{d}$ with samples $x\in\mathbb{R}^{d}$, flow-based models aim to learn a time-dependent velocity field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ leading to a flow $\phi:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that can push noise $x_{0}\sim p_{0}(x_{0})$ from source distribution to data $x_{1}\sim p_{1}(x_{1})$ from target distribution. This velocity field and associated flow can be defined by an ordinary differential equation (ODE): $\frac{d}{dt}\phi_{t}(x)=v_{t}(\phi_{t}(x),t)$ where $\phi_{0}(x)=x$. Similar to the denoising network in diffusion models, the Flow Matching (FM) objective trains a time-dependent network $v_{\theta}(x_{t},t)$ to regress against the ground truth velocity field $u_{t}(x_{t},t)$. However, direct computation of this FM objective is intractable in practice, since there is no closed-form solution of $u_{t}(x_{t},t)$. Instead, we can minimize the tractable Conditional Flow Matching (CFM) objective defined as:
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,p_{1}(x_{1}),p_{t}(x_{t}|x_{1})}\|v_{\theta}(x_{t},t)-u_{t}(x_{t},t|x_{1})\|^{2},$ (7)
where $t\sim\mathcal{U}(0,1)$, $x_{1}\sim p_{1}(x_{1})$, and $x_{t}\sim p_{t}(x_{t}|x_{1})$. It has been validated that the FM and CFM objectives share identical gradients with respect to $\theta$, while CFM offers the flexibility to choose the design choices of $u_{t}(x_{t}|x_{1})$ and $p_{t}(x_{t}|x_{1})$. A natural choice is to build the conditional probability paths as straight paths between the source and target distributions (Liu et al., 2023b; Lipman et al., 2023)i.e., $x_{t}=tx_{0}+(1-t)x_{1}$. We can then use Equation 7 for training and solve the ODE from $t=0$ to $t=1$ to sample new data points.
C.2 Classifier-free Guidance for Multi-Modal Conditions
Classifier-free guidance (Ho & Salimans, 2022) is a technique that balances quality and diversity in images generated by a diffusion model. For our PixWizard, the network $e_{\theta}(z_{t},c_{I},c_{T})$ has two conditionings: the input image $c_{I}$ and text instruction $c_{T}$. Similar to InstructPix2Pix (Brooks et al., 2023) and InstructCV (Gan et al., 2024), we employ a tailored noise predictor that assigns different weights, $w_{I}$ and $w_{T}$, to different conditionings, which can be adjusted to trade off how strongly the generated samples correspond with the input image and how strongly they correspond with the edit instruction. During training, we randomly set $c_{I}=\varnothing_{I}$ or $c_{T}=\varnothing_{T}$ for 5% of examples, and both conditions are $\varnothing$ for 5% of examples. The process is as follows:
$\begin{split}\tilde{e_{\theta}}(z_{t},c_{I},c_{T})=&\>e_{\theta}(z_{t},\varnothing,\varnothing)\\ &+w_{I}\cdot(e_{\theta}(z_{t},c_{I},\varnothing)-e_{\theta}(z_{t},\varnothing,\varnothing))\\ &+w_{T}\cdot(e_{\theta}(z_{t},c_{I},c_{T})-e_{\theta}(z_{t},c_{I},\varnothing))\end{split}$ (8)
C.3 Inference Efficiency with Fewer Steps
Compared to diffusion-based methods (Gan et al., 2024; Hu et al., 2024; Wang et al., 2023c), our model demonstrates better inference efficiency under a low number of function evaluations (NFE) regime. By leveraging the linear interpolation schedule with flow matching objective, PixWizard can learn straighter trajectories, thus requiring fewer sampling steps to achieve decent performance. Drawing inspiration from the observation that flow-based models exhibit higher discretization error at the beginning of generation, we employ the shifted time schedule which allocates more steps near noise. By further integrating high-order ODE solvers with improved schedules, such as Heun and Midpoint, our model produces reasonable output using only 30-60 NFEs, compared to 100 NFEs used in diffusion-based counterparts.
Appendix D More Experimental Results
D.1 Image Restoration
Following previous works (Conde et al., 2024; Potlapalli et al., 2024; Ai et al., 2024), besides the results from the deraining and denoising benchmarks, we selected six additional image restoration tasks to further evaluate the robustness of PixWizard: Snow100K-L (Liu et al., 2018) for desnowing, Reside (outdoor) SOTS (Li et al., 2018) for dehazing, LOLv2 (Yang et al., 2021) for low-light enhancement, and GoPro (Nah et al., 2017) for deblurring. Additionally, to assess zero-shot capabilities on tasks not encountered during training, we use TOLED (Zhou et al., 2021) for under-display camera (UDC) image restoration and UIEB (Li et al., 2019a) for underwater (UW) image restoration. We evaluate performance using PSNR and SSIM as distortion metrics.
Table 4: Quantitative results on 6 restoration tasks with existing image restoration methods.
Desnowing Dehazing Low-light Enh. Deblurring [Zero-shot](UDC)IR. [Zero-shot](UW)IR.
Methods Snow100K-L SOTS LOLv2 GoPro TOLED UIEB
PSNR↑ SSIM↑ PSNR↑ SSIM↑ PSNR↑ SSIM↑ PSNR↑ SSIM↑ PSNR↑ SSIM↑ PSNR↑ SSIM↑
SwinIR (2021) - - 21.50 0.891 - - 24.52 0.773 - - - -
Restormer (2022) 30.98 0.914 24.09 0.927 20.77 0.851 27.22 0.829 27.74 0.841 17.34 0.770
NAFNet (2022a) 31.42 0.920 25.23 0.939 18.04 0.827 26.53 0.808 27.90 0.848 17.31 0.736
AirNet (2022) 30.14 0.907 21.04 0.884 19.69 0.821 24.35 0.781 26.76 0.799 17.09 0.761
PromptIR (2024) 30.91 0.913 30.58 0.974 21.23 0.860 27.02 0.798 - - - -
DA-CLIP (2023) 28.31 0.862 30.16 0.936 21.76 0.762 24.65 0.703 - - - -
InstructIR (2024) - - 26.90 0.952 - - 29.70 0.892 - - - -
PixWizard 29.66 0.883 28.14 0.937 20.29 0.807 24.68 0.774 27.22 0.826 16.99 0.752
Results. Table 4 presents a additional performance comparison with state-of-the-art (SOTA) task-specific and all-in-one restoration methods. As shown in the results, even though image restoration data constitutes only a small portion of the overall training data, our method still demonstrates competitive performance, even outperforming some task-specific methods. For example, it outperforms DA-CLIP in the desnowing task and exceeds NAFNet, Restormer, and AirNet in the dehazing task. Furthermore, when evaluated on tasks not included in the training phase, our method achieved performance comparable to that of specialized models, highlighting the generalizability of our approach.
D.2 Image Grounding
We evaluate referring segmentation tasks on the gRefCOCO (gRef) (Liu et al., 2023a), RefCOCO, and RefCOCO+ validation and test sets. To assess the performance gap between our approach and specialized models, we report results from several expert methods and primarily compare our approach with two unified models: Unified-IO (Lu et al., 2022b) and InstructDiffusion (Geng et al., 2024). Unified-IO directly produces the corresponding binary mask, while InstructDiffusion requires a post-processing network to extract masks from the output image. We use two comparison methods: $(i)$ We convert the original image to the HSV color space to enhance the mask’s hue, then apply a threshold for extraction. These results are reported as w/ HSV. $(ii)$ We directly generate the corresponding binary mask for comparison. Following standard practices (Liu et al., 2023a), we use cumulative IoU (cIoU) to measure performance.
Table 5: Quantitative results on referring segmentation in terms of cIoU.
Method gRefCOCO RefCOCO RefCOCO+
val val test A test B val test A test B
LAVT (2022) 57.64 72.73 75.82 68.79 56.86 62.29 48.14
ReLA (2023a) 62.42 73.21 75.24 68.72 56.10 62.26 47.89
Unified-IO (2022a) 17.31 46.42 46.06 48.05 40.50 42.17 40.15
InstructDiffusion w/ HSV (2024) 33.19 41.64 40.81 41.98 33.20 37.85 26.92
PixWizard w/ HSV 33.65 47.28 44.12 45.38 40.07 38.89 40.76
PixWizard 34.17 46.44 41.25 43.43 36.49 35.30 38.93
Results. Table 5 reports the results for referring segmentation. Our model demonstrates strong performance under two different mask extraction methods, outperforming InstructDiffusion across almost all evaluation datasets. In the various test sets of RefCOCO and RefCOCO+, our performance is comparable to Unified-IO, but on gRefCOCO, we significantly outperform Unified-IO (34.17 vs. 17.31). However, it is important to note that these unified methods still lag considerably behind those specifically designed for referring segmentation. Additionally, we observed that using a generative model for grounding tasks presents a significant challenge due to the model’s limited perception and localization capabilities. This often results in failures to correctly locate objects when multiple targets are present, highlighting the need for improvements in instruction and understanding in future work.
D.3 Ablation Study
We performed ablation studies to assess the impact of each component’s design and the training process on learning in PixWizard. Given computational limitations, we conducted the ablation on PixWizard, training it for 40k steps.
Table 6: Comparison of the models with two different guidances, dynamic semantic tokens sampling (DSTS), and two-stage training and data balancing strategy for different tasks.
Deraining RefCOCO Depth Estim. Image Editing Canny-to-Image Inpainting
Methods Rain100L val NYUv2 Emu Edit MultiGen-20M Places
PSNR↑ cIoU↑ RMSE↓ $\text{CLIP}_{dir}\!\uparrow$ F1↑ FID↓ FID↓
M1 29.91(-0.33) 40.78(-0.94) 0.319(+0.005) 0.078(-0.010) 32.98(-0.03) 17.41(-0.27) 10.91(-0.03)
M2 14.72(-15.52) 18.43(-23.29) 0.586(-0.262) 0.071(-0.017) 11.12(-23.89) 19.34(-2.20) 13.87(-2.99)
PixWizard w/o DSTS 30.19(-0.05) 41.66(-0.06) 0.326(-0.002) 0.091(+0.003) 33.12(+0.11) 16.98(+0.16) 10.94(-0.06)
PixWizard w/o two-stage 29.17(-1.07) 40.18(-1.54) 0.322(+0.002) 0.085(-0.003) 33.07(+0.06) 17.21(-0.07) 11.07(-0.19)
PixWizard 30.24 41.72 0.324 0.088 33.01 17.14 10.88
Structural-Aware vs. Semantic-Aware Guidance. We demonstrate the importance of incorporating both structure-aware and semantic-aware guidance to boost PixWizard’s performance across a variety of heterogeneous tasks. To illustrate this, we trained two additional models: $(M1)$ using only the structure-aware module, and $(M2)$ using only the semantic-aware module. As shown in Table 6, the $M1$ model performs better on tasks that require preserving overall image structure and details. In contrast, the $M2$ model, which relies solely on cross-attention for injecting image features, struggles with most visual tasks, often generating images that deviate significantly from the input. However, as noted by other methods (Ye et al., 2023; Hu et al., 2024), semantic guidance typically excels in tasks that require more flexibility, such as text-to-image generation, conditional generation, and image editing. Compared to PixWizard, which integrates both modules, the combined approach enables the model to handle a wider variety of tasks, achieving balanced performance across different scenarios.
Influence of Dynamic Semantic Tokens Sampling. As shown in Table 6, the impact of using or not using Dynamic Semantic Tokens Sampling on task performance is minimal. This suggests that it’s unnecessary to model semantic tokens for the entire image. Since different tasks focus on different semantic areas, dynamically sampling the required features allows the model to handle each task more efficiently. Additionally, using fewer tokens reduces the computational load during attention calculations, resulting in faster inference times.
Influence of Two-Stage Training and Data Balancing. Table 6 presents the results of our proposed two-stage training and data balancing strategy. As shown, our approach is crucial, as the two-stage method significantly improves performance on tasks with smaller datasets while maintaining the same number of training steps and achieving faster convergence. Notably, for tasks with larger datasets, the performance remains comparable.